Two rigid rods as shown, AB and BC are connected to a pivot (A) and initially at rest as shown. Each rod has mass of 3 kg and length L = 0.5 m. A force P as shown is then applied to rod BC at some distance b from B (the pivot between bars A and B). Please find the distance b where, when force P is applied, the two rods (AB and BC) move together as if they formed a single rigid body of length 2L. (Hint: think about what it means for both bars to move together as a rigid body)
Answer #1

Dear student

Answer summary

b = 0.22827 m = 227.27 mm

Please pay attention to the kinetic diagram details.

All the best

Froee body diagram Enteral forces and internal reactions Kinetic dicqarum This diagram include accelcrcation toom (man) and rFree body diagrams um troms Kinetic clicgrom terms. ma and Ia at mass center Getermal force acting on body = Sinsilarly MomenFree body diagrams um troms Kinetic clicgrom terms. ma and Ia at mass center Getermal force acting on body = Sinsilarly Momenbefore calculating moment determine Mass, length, Qq and I total mass = m+m = 2m total lengthy= 24 MI of rod ML2 2mx (@L? abFBO of BC Kinetic diagram со всеми Cloud + veaction) (max+ I.) ł acts at CG Before starting calculation we have to find an (wtake moment about B Dx b = m Anex I + Ia - I = M2 au = 3 La Pb = mx 2 x x + 2 + M1L2 2 2 Pb = 3 Pb = q (Ma) intems of Ph yuz1-15b-b 녀b 야유 L = 0.5m b = 1x05 = 22721) b=227.27mm) Answer.